There are 12 movies showing at a theater. Four are action movies, two are children's movies and the other movies are comedies. Kevin randomly selects 2 different movies to see. He will not see a movie twice. What is the probability that the 1st movie is an action movie and the second movie is a comedy?


1
Expert's answer
2021-03-02T00:56:11-0500

"P=\frac{4}{12}*\frac{6}{11}=\frac{2}{11}."